A motor capable of producing a constant torque of 100Nm and maximum rotation speed of 150 rad/s is

connected to a flywheel with rotational inertia 0.1 kg-m2

. What angular acceleration will the flywheel

experience as the motor is switched on?


1
Expert's answer
2021-03-26T11:38:46-0400

Acording to the main equation of the rotational dynamics, obtain:


"\tau = I\varepsilon"

where "\tau = 100N\cdot m" is the applied torque, "I = 0.1kg\cdot m^2" is the moment of inertia, and "\varepsilon" is the angular acceleration. Thus, obtain:


"\varepsilon = \dfrac{\tau}{I}\\
\varepsilon = \dfrac{100N\cdot m}{0.1kg\cdot m^2} = 1000\space rad/s^2"

Answer. "1000\space rad/s^2".